OpenStudy (anonymous):

Simplify the expression. (c^3)^5(d^3)^0

OpenStudy (luckyness):

(c^3^)5(d^3)^0 =c^15

OpenStudy (algorithm):

zero right?

OpenStudy (mathstudent55):

You need to know these rules: \(\large (a^m)^n = a^{mn} \) \(\large a^0 = 1\)

OpenStudy (anonymous):

Thanks @Luckyness

OpenStudy (mathstudent55):

Start with your problem: \(\large (c^3)^5(d^3)^0\) Apply the first rule first: \(\large = c^{3 \times 5} d^{3 \times 0} \) \(\large =c^{15} d^0\) Now apply the second rule: \(\large =c^{15} \times 1 \) \(\large = c^{15} \)

OpenStudy (algorithm):

c to the power 15

OpenStudy (luckyness):

I agree with all of that! :P